# Leveraging Representations from Intermediate Encoder-blocks for Synthetic Image Detection

### v1

@@ 5.1 Comparative analysis
In Tables 2 and 3, we present the performance scores (ACC & AP respectively) of our method versus the competing ones. Our 1-class model outperforms all state-of-the-art methods irrespective of training class number. On average, we surpass the state-of-the-art by +9.4% ACC & +4.3% AP with the 1-class model, by +6.8% ACC & +4.4% AP with the 2-class model, and by +10.6% ACC & +4.5% AP with the 4-class model. In terms of ACC, we obtain the best score in 14 out of 20 test datasets, and simultaneously the first and second best performance in 10 of them. In terms of AP, we obtain the best score in 15 out of 20 test datasets, and simultaneously the first and second best performance in 14 of them. The biggest performance increase achieved by our method is on the SAN dataset [47] (+16.9% ACC).
@@
$\dagger$ Patch-Forensics has been trained on ProGAN data but not the on same dataset as the rest models. For more details please refer to [10].
@@
$\dagger$ Patch-Forensics has been trained on ProGAN data but not the on same dataset as the rest models. For more details please refer to [10].
@@ 5.2 Ablations
In Table 4, we present an ablation study, where we remove three of RINE’s main components, namely the intermediate representations, the TIE, and the contrastive loss, one-by-one. To be more precise, “w/o intermediate” means that we use only the last layer’s features (equivalent to the SotA method [16]). We measure the RINE’s performance on 20 test datasets, after removing each of the three components, and present the average ACC and AP for the 1-, 2-, and 4-class models, as well as the average across the three models. In addition, we present the respective results in terms of generalization, namely for the non-GAN generators only. The results demonstrate the positive impact of all proposed components, as the full architecture achieves the best performance. Also, ablating intermediate representations yields the biggest performance loss reducing the metrics to the previous state-of-the-art levels (i.e., [16]), as expected.
@@ 5.3 Hyperparameter analysis
Figure 3(a)-(c) illustrates ACC & AP boxplots for ViT-B/32 vs. ViT-L/14 in the 1-, 2-, and 4-class settings. Each boxplot is built from 48 scores (ACC/AP) obtained from the experiments of all combinations of $\xi$, $q$, and $d^{\prime}$. It is clear that CLIP ViT-L/14 outperforms ViT-B/32, thus from now on we report results of ViT-L/14 only. Figure 3(d)-(f) illustrates the impact of the contrastive loss factor $\xi$, showing little variability across the different values, in all training settings.
@@
Figure 4 presents ACC boxplots for each ($q$, $d^{\prime}$) combination of the trainable part of the model. For $q$=1, little variability is observed for different $d^{\prime}$ choices, while for $q$=2 and $q$=4, there is a slight performance increase with $d^{\prime}$ on average. The hyperparameter choices we have made for the proposed models are based on the maximum ACC & AP and not their average or distribution.
@@ 5.4 The effect of training duration
As stated in Section 4.2, the proposed models are trained for only one epoch. Here, we experimentally assess the potential benefit from further training. We identify the 3 best-performing configurations in the 1-epoch case for each of the 1-, 2-, and 4-class settings. Then, we train the corresponding models for 15 epochs and evaluate them in the 3rd, 5th, 10th, and 15th epoch (reducing learning rate by a factor of 10 at 6th and 11th epoch). Figure 5 presents their average and maximum performance, across the 20 test datasets, at different training stages. Neither the average nor the maximum performance improve if we continue training; instead, in the 1- and 4-class training settings performance drops.
@@ 5.5 The effect of training set size
Here, we present a set of experiments in order to assess the effect of the training set size to our models’ performance. Specifically, we re-train (again for only one epoch) the best models (1-class, 2-class, 4-class) using 20%, 50%, 80%, and 100% of the training data. For the 1-class, these percentages correspond to 7K, 18K, 28K, and 36K images, for the 2-class to 14K, 36K, 57K, and 72K images, and for the 4-class to 28K, 72K, 115K, and 144K images. In Figure 6, we present the obtained performance (ACC & AP) on the 20 test datasets. We observe that the results are very close to each other, especially in terms of AP, showcasing the effectiveness of the proposed method even in limited training data settings.
@@ 5.6 Robustness to common image perturbations
Following common practice [21, 9, 28, 16], in order to assess our method’s robustness to perturbations typically applied in social media, we assess our method’s performance after applying blurring, cropping, compression, noise addition, and their combination to each evaluation sample with 0.5 probability. The results in Table 5 indicate high robustness against cropping and compression, with minimal performance losses, and a good level of robustness against blurring and noise addition. Their combination seems to considerably reduce the models’ performance but at similar levels to that of current state-of-the-art without applying perturbations (cf. Table 2).
@@ 5.7 Importance of intermediate stages
In Section 5.2, it was demonstrated that the intermediate representations comprise the most important component of the proposed method as their incorporation produces the biggest performance increase. Here, we estimate the importance of each intermediate stage (i.e., Transformer block) by analysing the importance scores $\mathbf{A}=\{\alpha_{lk}\}$ estimated by the TIE module (cf. Equation 4) as follows:
@@
Figure 7 illustrates the frequency $f_{l}$ of Transformer block $l$ obtaining maximum importance, for $l$=1,…,$n$. For the 1-class and 4-class models $n$=24 and $d^{\prime}$=1024, while for the 2-class model $n$=24 and $d^{\prime}$=128. We observe that there are many Transformer blocks obtaining higher maximum importance frequency than the last one in all training settings, supporting the proposed method’s motivation.
@@ 5.8 The effect of training data origin on detecting Diffusion images
We additionally assess our method’s performance on Diffusion data produced by commercial tools, and more precisely on the Synthbuster dataset [56]. Due to the images’ size we consider the ten-crop validation technique and average the predictions. Table 6 illustrates RINE’s performance on Synthbuster when trained on ProGAN data from [9], and on Latent Diffusion Model (LDM) data from [15]. As expected training on Diffusion data provides big performance gains.
@@ Acknowledgments
This work is partially funded by the Horizon 2020 European projects vera.ai under grant agreement no. 101070093 and AI4Media under grant agreement no. 951911.

### v2

@@ 5.1 Comparative Analysis
In table˜2 and table˜3, we present the performance scores (ACC & AP respectively) of our method versus the competing ones. Our 1-class model outperforms all state-of-the-art methods irrespective of training class number. On average, we surpass the state-of-the-art by +9.4% ACC & +4.3% AP with the 1-class model, by +6.8% ACC & +4.4% AP with the 2-class model, and by +10.6% ACC & +4.5% AP with the 4-class model. In terms of ACC, we obtain the best score in 14 out of 20 test datasets, and simultaneously the first and second best performance in 10 of them. In terms of AP, we obtain the best score in 15 out of 20 test datasets, and simultaneously the first and second best performance in 14 of them. The biggest performance gain is on the SAN dataset [47] (+16.9% ACC). Also, considering more training classes does not reduce generalization (cf. Appendix).
@@
$\dagger$ Patch-Forensics has been trained on ProGAN data but not on the same dataset as the rest models. For more details please refer to [10].
@@
$\dagger$ Patch-Forensics has been trained on ProGAN data but not on the same dataset as the rest models. For more details please refer to [10].
@@ 5.2 Ablations
In table˜4, we present an ablation study, where we remove three of RINE’s main components, namely the intermediate representations, the TIE, and the contrastive loss, one-by-one. To be more precise, “w/o intermediate” means that we use only the last layer’s features (equivalent to the SotA method [16]). We measure the RINE’s performance on 20 test datasets, after removing each of the three components, and present the average ACC and AP for the 1-, 2-, and 4-class models, as well as the average across the three models. In addition, we present generalizability results by averaging across the non-GAN generators. The results demonstrate the positive impact of all proposed components. Also, ablating intermediate representations yields the biggest performance loss reducing the metrics to the previous SotA levels (i.e., [16]). In the Appendix the interested reader can find ablation of the fusion mechanism and backbone.
@@ 5.3 Hyperparameter Analysis
Figure˜3(a)-(c) illustrates ACC & AP boxplots for ViT-B/32 vs. ViT-L/14 in the 1-, 2-, and 4-class settings. Each boxplot is built from 48 scores (ACC/AP) obtained from the experiments of all combinations of $\xi$, $q$, and $d^{\prime}$. It is clear that CLIP ViT-L/14 outperforms ViT-B/32, thus from now on we report results of ViT-L/14 only. Figure˜3(d)-(f) illustrates the impact of the contrastive loss factor $\xi$, showing little variability across the different values, in all training settings.
@@
Figure˜4 presents ACC boxplots for each ($q$, $d^{\prime}$) combination of the trainable part of the model. For $q$=1, little variability is observed for different $d^{\prime}$ choices, while for $q$=2 and $q$=4, there is a slight performance increase with $d^{\prime}$ on average. The hyperparameter choices we have made for the proposed models are based on the maximum ACC & AP and not their average or distribution.
@@ 5.4 The Effect of Training Duration
As stated in section˜4.2, the proposed models are trained for only one epoch. Here, we experimentally assess the potential benefit from further training. We identify the 3 best-performing configurations in the 1-epoch case for each of the 1-, 2-, and 4-class settings. Then, we train the corresponding models for 15 epochs and evaluate them in the 3rd, 5th, 10th, and 15th epoch (reducing learning rate by a factor of 10 at 6th and 11th epoch). Figure˜5 presents their average and maximum performance, across the 20 test datasets, at different training stages. Neither the average nor the maximum performance improve if we continue training; instead, in the 1- and 4-class training settings performance drops.
@@ 5.5 The Effect of Training Set Size
Here, we present a set of experiments in order to assess the effect of the training set size to our models’ performance. Specifically, we re-train (again for only one epoch) the best models (1-class, 2-class, 4-class) using 20%, 50%, 80%, and 100% of the training data. For the 1-class, these percentages correspond to 7K, 18K, 28K, and 36K images, for the 2-class to 14K, 36K, 57K, and 72K images, and for the 4-class to 28K, 72K, 115K, and 144K images. In fig.˜6, we present the obtained performance (ACC & AP) on the 20 test datasets. We observe that the results are very close to each other, especially in terms of AP, showcasing the effectiveness of the proposed method even in limited training data settings.
@@ 5.6 Robustness to Common Image Perturbations
Following common practice [21, 9, 28, 16], we assess our method’s robustness to perturbations typically applied in social media by applying blurring, cropping, compression, noise addition, and their combination to each evaluation sample with 0.5 probability (experimental setting directly taken from [11]). The results in table˜5 indicate high robustness against cropping and compression, with minimal performance losses, and a good level of robustness against blurring and noise addition. Their combination considerably reduces performance but at similar levels to that of current SotA without applying perturbations (cf. table˜2).
@@ 5.7 Importance of Intermediate Stages
In section˜5.2, we show that the intermediate representations is the most important component of the proposed method. Here, we estimate the importance of each block by analysing the importances $\mathbf{A}=\{\alpha_{lk}\}$ estimated by TIE (cf. eq.˜4):
@@
[…]
@@ 5.8 The Effect of Training Data Origin on Detecting Diffusion Images
We additionally assess our method’s performance on Diffusion data produced by commercial tools, and more precisely on the Synthbuster dataset [56]. Due to the images’ size we consider the ten-crop validation technique and average the predictions. Table˜6 illustrates RINE’s performance on Synthbuster when trained on ProGAN data from [9], and on Latent Diffusion Model (LDM) data from [15]. As expected training on Diffusion data provides big performance gains.
Table 6: Performance (ACC/AP) of our method on Synthbuster.
@@ Acknowledgments
This work is partially funded by the Horizon Europe project vera.ai (GA No. 101070093) and H2020 project AI4Media (GA No. 951911).
Appendix
description GAN GAN GAN2 GAN GAN GAN GAN fake SITD SAN CRN IMLE Guided steps CFG steps 27 27 10 DALL-E
ACC
CLIP w/ FPN 99.1 81.8 75.0 90.0 86.1 68.4 99.3 55.1 72.8 50.7 52.6 60.7 53.1 77.6 54.8 78.9 58.2 58.4 57.6 61.9 69.6
RN50 w/ FPN 89.7 69.3 63.7 58.4 79.9 62.3 77.1 52.2 70.3 60.3 80.0 90.4 61.8 63.9 53.5 64.9 64.5 65.0 64.2 52.8 67.2
RINE 20-class 100.0 90.9 94.2 99.4 99.3 99.6 99.7 72.4 92.8 60.3 92.7 96.9 73.8 98.5 94.7 98.7 82.5 88.5 84.7 95.3 90.7
UFD 4-class 99.6 82.0 72.7 94.8 99.1 96.4 99.4 71.4 63.6 56.2 65.0 82.6 72.0 94.6 73.4 94.9 75.5 75.9 74.8 87.4 81.6
RINE w/ ViT 81.0 55.4 58.1 64.4 73.2 62.9 66.9 56.0 62.2 49.8 66.5 67.5 54.9 56.0 47.3 56.3 55.5 55.7 57.0 54.9 60.1
RINE w/ Wang 85.6 63.7 60.0 58.8 77.5 62.5 73.5 52.3 76.1 56.2 61.6 79.2 59.4 65.4 54.0 65.6 63.6 63.2 62.8 52.8 64.7
RINE 4-cl. (ours) 100.0 88.9 94.5 99.6 99.3 99.5 99.8 80.6 90.6 68.3 89.2 90.6 76.1 98.3 88.2 98.6 88.9 92.6 90.7 95.0 91.5
AP
CLIP w/ FPN 100.0 99.4 98.6 99.8 99.3 99.7 100.0 93.8 84.7 85.7 99.1 99.9 90.2 99.8 97.1 99.9 98.9 98.8 98.6 98.5 97.1
RN50 w/ FPN 96.8 79.6 73.6 65.1 85.5 87.3 88.1 56.8 82.2 70.8 95.2 98.1 69.6 73.1 55.8 74.1 70.9 72.5 71.7 54.2 76.0
RINE 20-class 100.0 99.8 100.0 100.0 100.0 100.0 100.0 97.5 97.8 84.2 98.1 99.9 95.4 99.9 99.3 99.9 96.7 98.1 96.9 99.4 98.1
UFD 4-class 100.0 96.7 98.7 99.3 99.9 99.7 100.0 85.5 64.0 76.7 94.6 99.0 88.0 99.4 92.3 99.3 94.1 94.5 93.8 97.7 93.6
RINE w/ ViT 91.0 57.6 64.0 71.4 83.3 71.8 76.2 59.4 77.1 52.2 68.8 72.1 56.9 57.5 45.7 58.3 55.1 55.7 56.4 54.5 64.3
RINE w/ Wang 94.5 76.8 73.6 66.3 85.1 87.1 85.6 58.2 80.3 60.0 87.2 96.3 65.9 77.9 56.7 77.4 76.3 77.5 75.6 57.0 75.8
RINE 4-cl. (ours) 100.0 99.4 100.0 99.9 100.0 100.0 100.0 97.9 97.2 94.9 97.3 99.7 96.4 99.8 98.3 99.9 98.8 99.3 98.9 99.3 98.8
Table 7: Further experimental results. Accuracy (ACC) and average precision (AP) are reported.
Feature fusion using Feature Pyramid Networks (FPN) [57]. Although RINE’s novelty lies in the use of intermediate CLIP features, rather than the (intentionally simple) fusion mechanism, we perform two experiments to assess the effectiveness of more complex fusion mechanisms such as that of Feature Pyramid Networks. In the first experiment (“CLIP w/ FPN” in Table 7 ) we replace $\mathfrak{Q}_{1}$, $\mathfrak{Q}_{2}$, and TIE modules by an FPN. In the second (“RN50 w/ FPN” in Table 7 ) we train a ResNet50 (pre-trained on ImageNet) with FPN. Incorporating FPN increases GPU memory consumption from 7GB to 28GB, and training time from 8min to 36min (1 epoch). Results are a lot worse than RINE in terms of ACC (69.6 vs. 91.5) and a little worse in terms of AP (97.1 vs. 98.8). Training of ResNet50 with FPN converges after 10 epochs. It consumes 5GB of GPU memory during training and needs 3min/epoch (31min in total). However, it still results in worse performance than RINE.
Fair comparison with UFD [16]. One could argue that the performance gain of RINE compared to UFD may be questionable, since training with more classes of ProGAN-generated images may make the model overfit to GAN-generated images and hurt its generalization capabilities. Thus, we additionally trained RINE on 20 classes (kept the 4-class configuration; performed no tuning), and UFD on 4 classes. The results are provided in Table 7 at lines “RINE 20-class” and “UFD 4-class”. RINE 20-class roughly preserves its 4-class instance performance (90.7 ACC, 98.1 AP) without any further tuning. No overfitting on GANs is observed. The frozen CLIP features are likely robust enough to prevent a lightweight MLP from overfitting. UFD maintains its performance in the 4-class setting as well, exhibiting no significant performance increase.
Considering simpler backbones. In order to assess the importance of the backbone choice (CLIP in RINE’s case) we perform two experiments using the ImageNet-pretrained ViT and Wang’s detector as backbones, respectively. Table 7 presents the results at lines “RINE w/ ViT” and “RINE w/ Wang”. The performance significantly decreases with these backbone choices.